# Periodic Cosmic String Formation and Dynamics

### v1

@@ 3.2 Vortex Formation Dynamics
As we saw in Sec. 2.4, parametric resonance drives the growth of modes with wavenumber $k\sim\mu$, eventually leading to the formation of counter-rotating regions. These regions of large $|\dot{\theta}|$ are separated by domain boundaries characterized by a vanishing angular speed, $\dot{\theta}\approx 0$, and large spatial gradients of the angular mode $\partial_{i}\theta$ (see also Fig. 6). This corresponds to the top row in Fig. 10. Prior to vortex formation, the location where $|\partial_{i}\theta|$ is large coincides with locations where $\rho/v$ is small, while the magnitude of the spatial component of the Noether current, $|j_{i}|=\rho^{2}|\partial_{i}\theta|$, is regular (in particular, non-zero) across the domain boundary. Note in general that the gradient field $\partial_{i}\theta$ is not divergence-free; i.e., $\partial^{i}\partial_{i}\theta\neq 0$. However, in the absence of a vortex, this gradient field is curl-free (by definition). Similar to the superfluid case RevModPhys.59.1001 , vortex production eventually occurs in the presence of the large spatial phase gradients at these domain boundaries (second row in Fig. 10). At these locations, vortex pair production may be facilitated by, for example, a passing field perturbation with a non-vanishing phase-gradient divergence. This perturbation can be seen explicitly in both the first and third columns of the first row in Fig. 10 (see also the animation at ZenodoVideos ). The interaction between this perturbation and the large-gradient domain boundary produces a pair of vortices (second row in Fig. 10). As a pair of vortices is formed, the radial mode and magnitude of the current density drop to zero at the production site, while the angular mode becomes multi-valued and $\partial_{i}\theta$ diverges at the vortex locations. The gradient field $\partial_{i}\theta$ swirls around the vortex core locations, implying non-vanishing vorticity [^18] Here, $\epsilon^{\mu\nu\lambda\rho}$ is the completely antisymmetric $(3+1)$-dimensional Minkowski spacetime Levi-Civita symbol with $\epsilon^{0123}\equiv+1$. In $(2+1)$ dimensions, we would simply drop the index $i$ in the expression in the text (2D spatial vorticity is a pseudoscalar); see also the definitions at footnote 19. $\epsilon^{0ijk}\partial_{j}\partial_{k}\theta$ at the center of the vortex core. Subsequently, this vortex–anti-vortex pair separates and begins propagating along the domain boundary separating the relevant counter-rotating regions (bottom row in Fig. 10). We will come back to the reason for this movement in Sec. 4. In the following subsections, we will first discuss the conditions for vortex formation in different regimes.
@@
Before we do, we pause to note that the vortices in Fig. 10 actually move superluminally (along the domain boundaries) at speeds $v\sim\mathcal{O}(3)$, at least for these parameter values; their speed would decrease for smaller $\mathcal{J}_{0}$ and for $\rho\approx v$, which is the limit in which familiar weakly coupled vortex results from condensed matter systems should hold. We stress that this superluminality is however not a violation of causality in our numerical simulations. Instead, this is a clear signal that one should not think of the vortices here as weakly coupled particle-like objects with a definite identity that exist independently of the background field in which they reside; rather, they are configurations that appear within the overall nonlinear causal evolution of the single field $\Phi$. We note that that this kind of apparent superluminality is actually familiar and happens even in completely prosaic linear wave-interference phenomena, if one fixes attention on specific field features instead of looking at the field evolution as a whole. For example, consider a pair of Gaussian line disturbances $f_{\pm}$ of amplitude $A$ and width $\sigma$, defined (in two dimensions) as
@@ 3.3.1 Small Initial Charge Densities $\mathcal{J}_{0}$
Consider the case where the initial charge density $\mathcal{J}_{0}$ is small, taking again as a concrete example the case we considered in Sec. 2.4.1: $D=2$, $\mathcal{J}_{0}=0.1\mu v^{2}$, $\rho_{i}/v=5$, and $\sigma_{\rho_{i}}=10^{-2}\rho_{i}$. The initial charge density is small here in the sense that the $\mathcal{J}_{0}^{2}/(2\rho^{2})$ term in the effective potential (13) is small compared to the other terms when evaluated at $\rho=\rho_{i}$, as well as in the sense that the magnitude of the typical angular-field temporal gradient $|\dot{\theta}|$ in the relevant counter-rotating regions immediately prior to vortex formation at their domain boundary is much larger than the magnitude of the initial global field rotation speed, $|\dot{\theta}_{i}|$. In this case, this initial charge density plays a limited role in the evolution of the system, and conditions for vortex formation can be found independent of the exact value of $\mathcal{J}_{0}$. We will discuss two conditions for vortex formation: a microscopic condition requiring a sufficiently large gradient field $\partial_{i}\theta$ at the location of vortex formation, and a global condition on the initial energy density in the system. We will also comment on the connections between these criteria.
In this example, counter-rotating regions form rapidly once the parametric resonance is active (see right panel of Fig. 5). In particular, the total spatial area of co- and counter-rotating regions (where the sense of rotation is defined with respect to the initial coherent state) is roughly the same during much of the exponential growth phase of the resonance (see the center panel of Fig. 6). Counter rotation persists through saturation (i.e., termination of the exponential growth caused by the parametric resonance) and vortex pair production. In Fig. 11, we show the state of the system shortly after the critical gradient $|\partial_{i}\theta|$ was surpassed along the domain boundaries, and (anti-)vortices have formed in the system. The field configuration in Fig. 11 is qualitatively representative of the post-saturation states of systems with $\mathcal{J}_{0}\lesssim v^{2}\mu$ and $\rho_{i}/v\gg 1$, when vortex formation actually occurs abundantly on domain boundaries whenever a critical gradient $|\partial_{i}\theta|$ is reached.
@@
The two critical conditions [i.e., sufficiently large $|\partial_{i}\theta|$, and (38)] are clearly related. As can be seen in Fig. 5, the energy that sustains counter rotation ($K_{\theta}$) arises via the parametric resonance instability from the the initial potential energy $V(\rho_{i})$. And a large $K_{\theta}$ (equivalently, a large $\dot{\theta}$ in the counter-rotating regions) is of course essential to drive the growth of the gradient $|\partial_{i}\theta|$ on the domain boundaries to the critical value for vortex formation. Moreover, as we will see in further examples below, with sufficient initial energy, counter-rotating regions persist even after vortices first form. This ensures that $|\partial_{i}\theta|$ can repeatedly reach the critical value as the difference of $\theta$ across the domain boundary periodically reaches $\Delta\theta\sim\mathcal{O}(2\pi)$, thereby causing repeated epochs of vortex formation.
@@ 3.3.2 Large Initial Charge Densities $\mathcal{J}_{0}$
Figure 12: Left panel: The critical initial radial field amplitude for vortex formation, $\rho_{c}$, as a function of $\mathcal{J}_{0}$, determined numerically here in $D=2$ with associated uncertainties [black points with errorbars]; see App. B for further details. This is confronted with the radial field amplitude $\rho_{\min}$ minimizing the effective radial-mode potential $V_{\rm eff}$ defined at (13) [blue line]. For large $\mathcal{J}_{0}$, the $\rho_{\min}$ scales as $\sim\mathcal{J}_{0}^{1/3}$, as shown [orange line]. Recall also that there is a second critical field value at $0<\rho_{c}^{(-)}<\rho_{\text{min}}$ which we do not show here; see discussion in footnote 21. Right panel: The difference between $\rho_{\min}$ and the critical field amplitude $\rho_{c}$. For large $\mathcal{J}_{0}$, this difference is roughly $\rho_{c}-\rho_{\min}\approx 0.13\ \mathcal{J}_{0}^{1/3}(v/\mu)^{1/3}$ [blue line]. We also indicate the value that the critical field amplitude takes in the $\mathcal{J}_{0}\rightarrow 0$ limit, $\rho_{c}/v\approx 1.24$ [horizontal gray line].
@@
In terms of conditions for vortex formation, these findings suggest that a large $\partial_{i}\theta$ is actually not sufficient for triggering vortex formation in regions where $\dot{\theta}$ is nonzero, and perhaps a Lorentz-invariant generalization of the condition that $|\partial_{i}\theta|$ be sufficiently large, is required. Specifically, motivated by the condition for Schwinger pair production in an electromagnetic field, we make an informed hypothesis for a generalization of this critical gradient $\partial_{i}\theta$ condition (cf. the critical velocity in a superfluid feynman1955chapter ): a critical value of $\kappa\equiv-|\partial_{\mu}\theta|^{2}\equiv(\partial_{i}\theta)^{2}-\dot{\theta}^{2}$, above which vortex pair production may occur. Such a generalization could be understood also from the perspective of the particle–vortex duality (see App. A and iengo1995vortex ), in which $\dot{\theta}$ maps onto a magnetic field232323 The $(2+1)$-dimensional Levi-Civita tensor $\epsilon$ is as defined in footnote 19.
@@
We can also find a minimum condition for vortex production on the initial field value $\rho_{i}$ for the large-$\mathcal{J}_{0}$ case (but, recall footnote 21). For each decade in $\mathcal{J}_{0}/(v^{2}\mu)$ between $10^{-1}$ and $10^{4}$, we determine the critical radial-field amplitude $\rho_{c}$, beyond which vortices form in our 2-dimensional simulations; see the right panel of Fig. 12. This critical initial field amplitude plateaus below $\mathcal{J}_{0}\sim v^{2}\mu$, but begins following a $\rho_{c}\propto\mathcal{J}_{0}^{1/3}$ trend towards large $\mathcal{J}_{0}$. Specifically, we find that a good fit to our data is provided by
@@
The place where $\mathcal{J}_{0}$ appears in this dimensionless equation is the dimensionless coefficient $\hat{\mu}^{2}\propto\mathcal{J}_{0}^{-2/3}$, which is subdominant in the large $\mathcal{J}_{0}$-limit. This suggests that, in this limit, the evolution of the system in terms of the fields $\hat{\rho}$ and $\hat{t}$ should be $\mathcal{J}_{0}$-independent during the growth of the perturbations; this in turn implies that the behavior of the system during this period should be almost invariant under scalings of the initial conditions that obey $\rho_{i}\propto\lambda^{-1/6}\mathcal{J}_{0}^{1/3}$. However, at the onset of the formation of the counter-rotating regions, the domain boundaries emerge as regions where $\dot{\theta}\approx 0$, implying both that $j_{0}$ [cf. (3)] becomes small on the boundaries, and also that there is a large field perturbation at the boundary location, which invalidates the naive application of the perturbative approach that led to (12) and thence to (41) in the understanding of the nonlinear field dynamics. Moreover, to the extent that (12) or (41) do still give some guidance, a term morally of the form $-\hat{\mu}^{2}\hat{\rho}_{0}$ will become relevant on the boundaries, invalidating the scaling we found above at large $\mathcal{J}_{0}$. Unfortunately, the locations where these failures occur are precisely the domain boundaries, which is where vortices form in pairs. As a result of all of these issues, we cannot prove analytically that $\rho_{c}-\rho_{\min}$ should scale as $\lambda^{-1/6}\mathcal{J}_{0}^{1/3}$. Nevertheless, the above argument is suggestive that we should, in the large $\mathcal{J}_{0}$ limit, expect similar system behavior for the perturbations as they grow toward the counter-rotating regions that precede vortex formation under the rescaling $\rho_{i}\propto\lambda^{-1/6}\mathcal{J}_{0}^{1/3}$ as $\mathcal{J}_{0}$ is varied. It is therefore not unexpected that similar vortex formation dynamics would be at play under such rescaling. And, indeed, the numerical results in fact indicate that the scaling $\rho_{c}-\rho_{\min}\propto\mathcal{J}_{0}^{1/3}$ does obtain in the large $\mathcal{J}_{0}$ limit.
@@ 3.4 Vortex Formation in Three Spatial Dimensions
In two spatial dimensions, pairs of vortices are efficiently produced on domain boundaries, where the magnitude of the spacetime gradient of the angular mode is sufficiently large. In three spatial dimensions, on the other hand, strings in our setting must form in loops.242424 By contrast, in condensed matter systems, strings may end on material boundaries tinkham2004introduction ; Galaiko1966FormationOV . Therefore, it is important that we verify the string-production mechanism in three spatial dimensions behaves similarly to its 2D counterpart. To that end, we numerically solve the full nonlinear scalar field evolution, determined by (4), in $D=3$ spatial dimensions. As before, we impose periodic boundary conditions and set initial conditions as described in Sec. 2.4. We focus on the example case $\mathcal{J}_{0}=0.1\mu v^{2}$, $\rho_{i}/v=5$, and $\sigma_{\rho_{i}}=10^{-2}\rho_{i}$.

### v2

@@ 3.2 Vortex Formation Dynamics
As we saw in Sec. 2.4, parametric resonance drives the growth of modes with wavenumber $k\sim\mu$, eventually leading to the formation of counter-rotating regions. These regions of large $|\dot{\theta}|$ are separated by domain boundaries characterized by a vanishing angular speed, $\dot{\theta}\approx 0$, and large spatial gradients of the angular mode $\partial_{i}\theta$ (see also Fig. 6). This corresponds to the top row in Fig. 10. Prior to vortex formation, the location where $|\partial_{i}\theta|$ is large coincides with locations where $\rho/v$ is small, while the magnitude of the spatial component of the Noether current, $|j_{i}|=\rho^{2}|\partial_{i}\theta|$, is regular (in particular, non-zero) across the domain boundary. Note in general that the gradient field $\partial_{i}\theta$ is not divergence-free; i.e., $\partial^{i}\partial_{i}\theta\neq 0$. However, in the absence of a vortex, this gradient field is curl-free (by definition). Similar to the superfluid case RevModPhys.59.1001 , vortex production eventually occurs in the presence of the large spatial phase gradients at these domain boundaries (second row in Fig. 10). At these locations, vortex pair production may be facilitated by, for example, a passing field perturbation with a non-vanishing phase-gradient divergence. This perturbation can be seen explicitly in both the first and third columns of the first row in Fig. 10 (see also the animation at ZenodoVideos ). The interaction between this perturbation and the large-gradient domain boundary produces a pair of vortices (second row in Fig. 10). As a pair of vortices is formed, the radial mode and magnitude of the current density drop to zero at the production site, while the angular mode becomes multi-valued and $\partial_{i}\theta$ diverges at the vortex locations. The gradient field $\partial_{i}\theta$ swirls around the vortex core locations, implying non-vanishing vorticity [^19] Here, $\epsilon^{\mu\nu\lambda\rho}$ is the completely antisymmetric $(3+1)$-dimensional Minkowski spacetime Levi-Civita symbol with $\epsilon^{0123}\equiv+1$. In $(2+1)$ dimensions, we would simply drop the index $i$ in the expression in the text (2D spatial vorticity is a pseudoscalar); see also the definitions at footnote 21. $\epsilon^{0ijk}\partial_{j}\partial_{k}\theta$ at the center of the vortex core. Subsequently, this vortex–anti-vortex pair separates and begins propagating along the domain boundary separating the relevant counter-rotating regions (bottom row in Fig. 10). We will come back to the reason for this movement in Sec. 4. In the following subsections, we will first discuss the conditions for vortex formation in different regimes.
@@
Before we do, we pause to note that the vortices in Fig. 10 actually move superluminally (along the domain boundaries) at speeds $v\sim\mathcal{O}(3)$, at least for these parameter values; their speed would decrease for smaller $\mathcal{J}_{0}$ and for $\rho\approx v$, which is the limit in which familiar weakly coupled vortex results from condensed matter systems should hold. We stress that this superluminality is however not a violation of causality in our numerical simulations. Instead, this is a clear signal that one should not think of the vortices here as weakly coupled particle-like objects with a definite identity that exist independently of the background field in which they reside; rather, they are configurations that appear within the overall nonlinear causal evolution of the single field $\Phi$. We note that this kind of apparent superluminality is actually familiar and happens even in completely prosaic linear wave-interference phenomena, if one fixes attention on specific field features instead of looking at the field evolution as a whole. For example, consider a pair of Gaussian line disturbances $f_{\pm}$ of amplitude $A$ and width $\sigma$, defined (in two dimensions) as
@@ 3.3.1 Small Initial Charge Densities $\mathcal{J}_{0}$
Consider the case where the initial charge density $\mathcal{J}_{0}$ is small, taking again as a concrete example the case we considered in Sec. 2.4.1: [^20] Recall, $\sigma_{\rho_{i}}$ sets the initial amplitude of the unstable modes; the production time of the first vortices depends logarithmically on $\sigma_{\rho_{i}}$. $D=2$, $\mathcal{J}_{0}=0.1\mu v^{2}$, $\rho_{i}/v=5$, and $\sigma_{\rho_{i}}=10^{-2}\rho_{i}$. The initial charge density is small here in the sense that the $\mathcal{J}_{0}^{2}/(2\rho^{2})$ term in the effective potential (13) is small compared to the other terms when evaluated at $\rho=\rho_{i}$, as well as in the sense that the magnitude of the typical angular-field temporal gradient $|\dot{\theta}|$ in the relevant counter-rotating regions immediately prior to vortex formation at their domain boundary is much larger than the magnitude of the initial global field rotation speed, $|\dot{\theta}_{i}|$. In this case, this initial charge density plays a limited role in the evolution of the system, and conditions for vortex formation can be found independent of the exact value of $\mathcal{J}_{0}$. We will discuss two conditions for vortex formation: a microscopic condition requiring a sufficiently large gradient field $\partial_{i}\theta$ at the location of vortex formation, and a global condition on the initial energy density in the system. We will also comment on the connections between these criteria.
In this example, counter-rotating regions form rapidly once the parametric resonance is active (see right panel of Fig. 5). In particular, the total spatial area of co- and counter-rotating regions (where the sense of rotation is defined with respect to the initial coherent state) is roughly the same during much of the exponential growth phase of the resonance (see the center panel of Fig. 6). Counter rotation persists through saturation (i.e., termination of the exponential growth caused by the parametric resonance) and vortex pair production. In Fig. 11, we show the state of the system shortly after the critical gradient $|\partial_{i}\theta|$ was surpassedalong the domain boundaries, and (anti-)vortices have formed in the system. This critical gradient that is reached locally on the domain boundaries prior to vortex formation is parametrically of order $|\partial_{i}\theta|\sim 2\pi/m^{\textsc{crr}}_{\rho}$, where $m^{\textsc{crr}}_{\rho}$ is the mass of the radial mode in the immediately adjacent CRR [parametrically, $m^{\textsc{crr}}_{\rho}\sim\mathcal{O}(\mu)$]. The field configuration in Fig. 11 is qualitatively representative of the post-saturation states of systems with $\mathcal{J}_{0}\lesssim v^{2}\mu$ and $\rho_{i}/v\gg 1$, when vortex formation actually occurs abundantly on domain boundaries whenever a critical gradient $|\partial_{i}\theta|$ is reached.
@@
Figure 12: Left panel: The critical initial radial field amplitude for vortex formation, $\rho_{c}$, as a function of $\mathcal{J}_{0}$, determined numerically here in $D=2$ with associated uncertainties [black points with errorbars]; see App. B for further details. This is confronted with the radial field amplitude $\rho_{\min}$ minimizing the effective radial-mode potential $V_{\rm eff}$ defined at (13) [blue line]. For large $\mathcal{J}_{0}$, the $\rho_{\min}$ scales as $\sim\mathcal{J}_{0}^{1/3}$, as shown [orange line]. Recall also that there is a second critical field value at $0<\rho_{c}^{(-)}<\rho_{\text{min}}$ which we do not show here; see discussion in footnote 23. Right panel: The difference between $\rho_{\min}$ and the critical field amplitude $\rho_{c}$. For large $\mathcal{J}_{0}$, this difference is roughly $\rho_{c}-\rho_{\min}\approx 0.13\ \mathcal{J}_{0}^{1/3}(v/\mu)^{1/3}$ [blue line]. We also indicate the value that the critical field amplitude takes in the $\mathcal{J}_{0}\rightarrow 0$ limit, $\rho_{c}/v\approx 1.24$ [horizontal gray line].
@@
The two critical conditions [i.e., sufficiently large $|\partial_{i}\theta|$, and (38)] are clearly related. As can be seen in Fig. 5, the energy that sustains counter rotation ($K_{\theta}$) arises via the parametric resonance instability from the initial potential energy $V(\rho_{i})$. And a large $K_{\theta}$ (equivalently, a large $\dot{\theta}$ in the counter-rotating regions) is of course essential to drive the growth of the gradient $|\partial_{i}\theta|$ on the domain boundaries to the critical value for vortex formation. Moreover, as we will see in further examples below, with sufficient initial energy, counter-rotating regions persist even after vortices first form. This ensures that $|\partial_{i}\theta|$ can repeatedly reach the critical value as the difference of $\theta$ across the domain boundary periodically reaches $\Delta\theta\sim\mathcal{O}(2\pi)$, thereby causing repeated epochs of vortex formation.
@@ 3.3.2 Large Initial Charge Densities $\mathcal{J}_{0}$
In terms of conditions for vortex formation, these findings suggest that a large $\partial_{i}\theta$ is actually not sufficient for triggering vortex formation in regions where $\dot{\theta}$ is nonzero, and perhaps a Lorentz-invariant generalization of the condition that $|\partial_{i}\theta|$ be sufficiently large, is required. Specifically, motivated by the condition for Schwinger pair production in an electromagnetic field, we make an informed hypothesis for a generalization of this critical gradient $\partial_{i}\theta$ condition (cf. the critical velocity in a superfluid feynman1955chapter ): a critical value of $\kappa\equiv-|\partial_{\mu}\theta|^{2}\equiv(\partial_{i}\theta)^{2}-\dot{\theta}^{2}$, above which vortex pair production may occur. Such a generalization could be understood also from the perspective of the particle–vortex duality (see App. A and iengo1995vortex ), in which $\dot{\theta}$ maps onto a magnetic field252525 The $(2+1)$-dimensional Levi-Civita tensor $\epsilon$ is as defined in footnote 21.
@@
We can also find a minimum condition for vortex production on the initial field value $\rho_{i}$ for the large-$\mathcal{J}_{0}$ case (but, recall footnote 23). For each decade in $\mathcal{J}_{0}/(v^{2}\mu)$ between $10^{-1}$ and $10^{4}$, we determine the critical radial-field amplitude $\rho_{c}$, beyond which vortices form in our 2-dimensional simulations; see the right panel of Fig. 12. This critical initial field amplitude plateaus below $\mathcal{J}_{0}\sim v^{2}\mu$, but begins following a $\rho_{c}\propto\mathcal{J}_{0}^{1/3}$ trend towards large $\mathcal{J}_{0}$. Specifically, we find that a good fit to our data is provided by
@@
The place where $\mathcal{J}_{0}$ appears in this dimensionless equation is the dimensionless coefficient $\hat{\mu}^{2}\propto\mathcal{J}_{0}^{-2/3}$, which is subdominant in the large $\mathcal{J}_{0}$-limit. This suggests that, in this limit, the evolution of the system in terms of the fields $\hat{\rho}$ and $\hat{t}$ should be $\mathcal{J}_{0}$-independent during the growth of the perturbations; this in turn implies that the behavior of the system during this period should be almost invariant under scalings of the initial conditions that obey $\rho_{i}\propto\lambda^{-1/6}\mathcal{J}_{0}^{1/3}$. However, at the onset of the formation of the counter-rotating regions, the domain boundaries emerge as regions where $\dot{\theta}\approx 0$, implying both that $j_{0}$ [cf. (3)] becomes small on the boundaries, and also that there is a large field perturbation at the boundary location, which invalidates the naive application of the perturbative approach that led to (12) and thence to (41) in the understanding of the nonlinear field dynamics. Moreover, to the extent that (12) or (41) do still give some guidance, a term morally of the form $\hat{\mu}^{2}\hat{\rho}_{0}$ will become relevant on the boundaries, invalidating the scaling we found above at large $\mathcal{J}_{0}$. Unfortunately, the locations where these failures occur are precisely the domain boundaries, which is where vortices form in pairs. As a result of all of these issues, we cannot prove analytically that $\rho_{c}-\rho_{\min}$ should scale as $\lambda^{-1/6}\mathcal{J}_{0}^{1/3}$. Nevertheless, the above argument is suggestive that we should, in the large $\mathcal{J}_{0}$ limit, expect similar system behavior for the perturbations as they grow toward the counter-rotating regions that precede vortex formation under the rescaling $\rho_{i}\propto\lambda^{-1/6}\mathcal{J}_{0}^{1/3}$ as $\mathcal{J}_{0}$ is varied. It is therefore not unexpected that similar vortex formation dynamics would be at play under such rescaling. And, indeed, the numerical results in fact indicate that the scaling $\rho_{c}-\rho_{\min}\propto\mathcal{J}_{0}^{1/3}$ does obtain in the large $\mathcal{J}_{0}$ limit.
@@ 3.4 Vortex Formation in Three Spatial Dimensions
In two spatial dimensions, pairs of vortices are efficiently produced on domain boundaries, where the magnitude of the spacetime gradient of the angular mode is sufficiently large. In three spatial dimensions, on the other hand, strings in our setting must form in loops.262626 By contrast, in condensed matter systems, strings may end on material boundaries tinkham2004introduction ; Galaiko1966FormationOV . Therefore, it is important that we verify the string-production mechanism in three spatial dimensions behaves similarly to its 2D counterpart. To that end, we numerically solve the full nonlinear scalar field evolution, determined by (4), in $D=3$ spatial dimensions. As before, we impose periodic boundary conditions and set initial conditions as described in Sec. 2.4. We focus on the example case $\mathcal{J}_{0}=0.1\mu v^{2}$, $\rho_{i}/v=5$, and $\sigma_{\rho_{i}}=10^{-2}\rho_{i}$.